# Towards Physical Understanding in Video Generation: A 3D Point Regularization Approach

### v1

@@ 3.2 Point Augmentation
With our 3D-aware video dataset, we now aim to extend the capabilities of the video model to generate not only the RGB video $\mathcal{V}$ but also the corresponding 3D representation $\mathcal{P}$. Aligning the video with the point cloud enables the model to inherently acquire 3D knowledge. The augmented video-point model learns not only how the object appears in a 2D color image but also its 3D geometry and location.
@@
Our augmentation pipeline is visualized in Figure 2. It is worth noting that training the video-point model entirely from scratch is unnecessary, as training a video diffusion model already demands substantial computational resources and vast amounts of video data. Therefore, instead of training from scratch, it is more practical to fine-tune an existing video model to incorporate the point-tracking modality. In this work, we employ a similar architecture to the video model in I2VGen-XL [41, 35], a UNet-based latent diffusion model conditioned on image and text prompts.
@@
To collectively learn from and generate $\mathcal{V}$ and $\mathcal{P}$, the straightforward way would be to feed them together to the diffusion model. Since we are matching the shapes of video and point tensors, we can concatenate them along the channel dimension, resulting in an point-augmented video $\mathcal{VP}$ with a shape of $[T,H,W,2C]$. This concatnated video can be seamlessly integrated into the UNet by simply adjusting the input/output dimensions to accommodate the additional channels. This increase in dimensionality does not overshadow the weights of the original model; instead, only new dimensions are introduced to handle the added modality. To leverage pretrained RGB weights, we can load the pretrained model for the corresponding RGB channels and initialize the remaining weights as zeros. For the conditioning frame, we continue using the first RGB frame and repeat the condition image tensor to prompt the generation of both video and point tracking, as our goal is for the model not to rely on the initial point positions during the inference stage. This strategy is effective because the same reference frame is used for generating the point-tracking data earlier.
@@
In addition to the pixel-wise alignment of video and point data in creating our dataset, we introduce a cross-attention layer at the entry of the UNet. This layer connects the first and second halves of the channel dimensions, corresponding to color information and spatial location in the latent dimension, to further enhance the alignment between them.
@@
Incorporating the aforementioned changes, The training pipeline then proceed as usual: we encode $\mathcal{VP}$ to latent space $z$, noise it to $z_{t}$ with randomly sampled noise $\epsilon$ and diffusion step $t$. The UNet model then predict the added noise and compare with true noise, where we adopt DDIM scheduling for the model. With the modified architecture, we fine-tune the diffusion model on PointVid to obtain our video-point generation model. Given an RGB image and a text prompt, our model can generate both the video and the corresponding 3D trajectories.
@@ 3.3 Point Regularization
With the video closely aligned to 3D trajectories, we leverage this additional modality to enhance RGB video generation. To achieve this, we introduce regularization terms on the point tracking component of the diffusion output, aligning it with the ground truth, which in turn implicitly improves the RGB video output. (see Figure 2).
@@
The first component we need is to generate directly usable point trackings for evaluation in 3D space. Since including random noise of any level in the point cloud would likely render it unusable, we employ DDIM sampling to recover a noise-free stage before applying regularization. In each training step with a latent input $z$ and corresponding noisy stage $z_{t}$, We apply DDIM denoising to recover $z_{0}$ iteratively. Following the approach of [40], we utilize gradient checkpointing to freeze the gradient flow except during the final step, preventing memory overflow. The denoised data $z_{0}$ is subsequently decoded into explicit 3D point tracking $\tilde{\mathcal{P}}$ as the sampled 3D trajectory of the video.
@@
Figure 4: Point Regularization. The reconstructed point cloud in the diffusion output often contains noise and deformations (middle). This issue is mitigated using our point regularization (right). The synthetic point cloud above (e.g. box and shoes falling on the ground) is generated by Kubric  [13] and trained with our pipeline.
@@
With the explicit modeling of trajectories as a 3D representation, we introduce regularization terms to improve the quality of the generated results, focusing on motion fidelity and temporal smoothness. For simplicity, let us assume $\tilde{\mathcal{P}}$ is reshaped to $T\times N\times 3$, and we use $\tilde{\mathcal{P}}^{\tau}$ to denote the point cloud of shape $N\times 3$ at time $\tau$.
@@
A common issue that arises during point tracking generation is the presence of excessive noise (see Figure 4 top middle). This noise can originate from various sources, such as the variational autoencoder (VAE) used to encode $\mathcal{P}$ into the latent space. High-frequency noise in both spatial and temporal dimensions tends to obscure the main shape and motion, thereby impairing the video model’s ability to track 3D points effectively. To mitigate noise and increase fidelity of reconstruction, we propose the reconstruction loss $\mathcal{L}_{\text{recon}}$ as follows:
$\mathcal{L}_{\text{recon}}=c_{0}\sum_{\tau=0}^{T}\|\tilde{\mathcal{P}}^{\tau}-\mathcal{P}^{\tau}\|+c_{1}\sum_{\tau=1}^{T}\|\tilde{\mathcal{P}}^{\tau}-\tilde{\mathcal{P}}^{\tau-1}\|\\ +c_{2}\sum_{\tau=2}^{T}\|\tilde{\mathcal{P}}^{\tau}-2\tilde{\mathcal{P}}^{\tau-1}+\tilde{\mathcal{P}}^{\tau-2}\|.$ (1)
@@
Here the first term penalizes the difference between generated point cloud and the noise-free ground truth $\mathcal{P}$, ensuring generation fidelity, while the second term penalizes the difference between consecutive time step to encourage temporal smoothness; the third term further enforeces acceleration smoothness, discouraging abrupt changes in velocity. We selected the weights so that each term are about the same scale initially.
@@
By combining reconstruction and rigidity loss, we enhance the generation quality of the 3D point cloud, as shown in Figure 4 (right). This improvement in the point space implicitly promotes shape preservation and smooth motion transitions in the RGB video, as our video-point model closely aligns the two modalities. To prevent the geometric regularizations from dominating the semantic aspects, we jointly optimize these regularization terms with diffusion losses during training, resulting in an overall loss of the form:
$\mathcal{L}=\lambda_{\text{ddim}}\mathcal{L}_{\text{ddim}}+\lambda_{\text{recon}}\mathcal{L}_{\text{recon}}+\lambda_{\text{rigid}}\mathcal{L}_{\text{rigid}}$ (3)
Figure 5: Ablation on Point Augmentation and Regularization. Our point-augmented model demonstrates a degree of 3D-awareness compared to the model fine-tuned on video data alone, but it still exhibits some artifacts, which are mitigated through regularization.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of our proposed method through a series of experiments. We collect around 70K videos from [9] and our proprietary dataset to form our training dataset. The videos are processed to a resolution of 448$\times$256 with 16 frames. During the augmentation stage, the video-point diffusion model is trained on 8 A100 GPUs with a batch size of 4. In the subsequent regularization stage, we train on a single A100 GPU with a batch size of 4. We evaluate the results both qualitatively and quantitatively, showcasing the model’s successful handling of contact-rich scenarios and its strong shape/motion consistency.
@@ 4.1 Qualitative Results
Although our model is designed to generate sequences of 2D images, our point augmentation and regularization enhance its ability to handle inherently 3D motions. To evaluate, we test our model on unseen images and visualize the corresponding RGB video. For comparison, we apply the same prompts to the baseline model I2VGen-XL [41], as well as two additional models: SVD (stable-video-diffusion-img2vid-xt) [6] and DynamiCrafter [37]. For qualitative comparison, we focus primarily on the aspects of object interaction and shape conservation.
Task-oriented videos
A prevalent category of scenarios that make up a significant portion of real-world videos in the wild involves humans performing specific tasks (see Figure 1), such as playing with a toy, cooking food, or handling tools. However, current video models often struggle to capture localized details, particularly how human hands interact with objects. The challenge arises from the complex 3D nature of the interacting parts, and relying solely on 2D data fails to capture how different parts shape, move, and contact. As a result, hands or objects may undergo severe distortion during interactions, leading to unrealistic artifacts. With our 3D-aware generation framework, our model effectively preserves the shape of both hands and objects, ensuring smooth transitions throughout the video, exhibiting consistent quality improvement over the compared models.
General videos
The strength of our model extends to generating high-fidelity videos across a variety of categories. In Figure 6, we showcase several categories: (a) dynamic objects, (b) static objects with a moving camera, (c/d) humans, and (e/f) animals. For objects with camera motion, our 3D-aware regularization enhances the model’s understanding of 3D shape and scene composition, effectively eliminating artifacts such as foreground object morphing and background degradation. For human subjects, our point augmentation prevents unnatural changes in shape and appearance. Additionally, our model accurately captures the shape and behavior of animals, avoiding morphing issues.
@@ 4.2 Quantitative Results
We quantitatively evaluate our model using batch inference on the test set. We apply the metrics to a test batch obtained from [9], consisting of a total of 387 video clips. We compare the scores with the base model, I2VGen-XL, and summarize the results in Table 1.
First, we assess the general aspects of video generation quality using VBench [16], evaluating the content and aesthetic quality of the videos. The results show that our method outperforms in most categories. Notably, our point regularization contributes to significant improvement in content consistency.
Beyond the general aspects, we focus on how 3D-awareness enhances alignment with real-world physical principles. It is important to note that determining whether a 2D video is physically correct is largely subjective, as ground-truth 3D physical states are not available. Therefore, we use VideoPhy [4], where the authors utilize human annotations to train a discriminator that provides a physical commonsense score. VideoPhy defines physical commonsense based on various perspectives, such as solid mechanics, where solid objects are expected to retain their shape and size throughout the video. As shown in Table 1, our model achieves a higher score for alignment with physical principles.
Figure 6: Comparison on General Videos. We showcase the generated videos across various categories, including static and dynamic objects, humans, and animals. Our method ensures smooth transitions in object shape and motion and eliminates morphing artifacts.
@@ 4.3 Ablation Study
We conduct ablation studies to demonstrate the essential role of each component of our method, namely point augmentation and regularization (see Figure 5). As a baseline, we finetune the diffusion model on our dataset using only video data to eliminate bias from training videos (Figure 5 left). Next, we perform inference using the jointly trained video-point model to highlight the improvement from point augmentation alone (Figure 5 middle). Finally, we evaluate the full model with regularization (Figure 5 right). We observe that the model trained exclusively with RGB video can generate videos that match the appearance and semantics of the prompts. However, it suffers from severe distortions in both hands and objects. By incorporating point tracking during training, we leverage the 3D information from PointVid to gain partial awareness of spatial locations. This enhancement is evident as objects in the scene (e.g., cellphone and white stars) maintain their correct 3D shapes. This partial awareness is further enhanced by our regularization steps, leading to better control of shape and motion, significantly reducing artifacts in the videos (e.g., hands now retain their shape).
@@ 5 Conclusion and Discussion
To summarize, we propose a generation framework that incorporates 3D-awareness into video diffusion models. By tracking objects in 3D and aligning them in pixel space, we elevate traditional video datasets to a new dimension, revealing out-of-plane information previously inaccessible to video models. Through joint training, the video model learns to perceive 3D shape and motion, acquiring physical commonsense that is natural in 3D. We then apply regularization to refine the generation process and further enhance the results, eliminating artifacts such as object morphing. Our method has demonstrated a strong capability to handle content-rich scenarios, such as task-oriented videos, and shows superior visual plausibility when generating videos of general categories.
@@
We acknowledge a few limitations of our framework. The degree of 3D-awareness, rooted in our video-point joint training, is constrained by the resolution of our 3D points. Objects that are not sufficiently represented may lack a quality 3D prior and may still exhibit implausible artifacts. Additionally, the generation quality of our model is somewhat limited by the backend UNet model, and could potentially be improved by employing more advanced transformer-based models. We leave these aspects as directions for future exploration.
Supplementary Material
Supplementary to the main paper, we provide more details on dataset creation in Appendix A, additional qualitative results in Appendix B, additional results for ablation studies in Appendix C, and a human preference evaluation in Appendix D.
Appendix A PointVid Dataset Generation
In this section, we present additional details on generating our 3D-aware video dataset, PointVid. A visualization of the generation pipeline is provided in the main paper. Our pipeline requires no additional annotations beyond the video-text pairs in the dataset. For our experiments, we utilize Panda70M [9].

### v2

@@ 3.2 Point Augmentation in Diffusion Denoiser
With our 3D-aware video dataset, we now aim to extend the capabilities of the video model to generate not only the RGB video $\mathcal{V}$ but also the corresponding 3D representation $\mathcal{P}$. Aligning the video with the point cloud enables the model to inherently acquire 3D knowledge. The augmented video-point model learns not only how the object appears in a 2D color image but also its 3D geometry.
@@
Our augmentation pipeline is illustrated in Figure 2. It is worth noting that training the video-point model entirely from scratch is unnecessary, as training video diffusion models already requires significant computational resources and large-scale datasets. Instead, it is more practical to fine-tune a pretrained video model to incorporate the point-tracking modality. Additionally, our injection of 3D points is agnostic to the network architecture of the pretrained video model. In this work, we employ both UNet- and DiT-based denoisers—specifically, I2VGen-XL zhang2023i2vgen and CogVideoX-1.5 yang2024cogvideox —both conditioned on text-image prompts. While our training pipeline incorporates an additional point modality, inference remains conditioned only on text-image pairs, consistent with the base model.
@@
To collectively learn from and generate $\mathcal{V}$ and $\mathcal{P}$, the straightforward way would be to feed them together to the diffusion model. Since we are matching the shapes of video and point tensors, we can concatenate them along the channel dimension, resulting in an point-augmented video $\mathcal{VP}$ with a shape of $[T,H,W,2C]$. This concatenated video can be seamlessly integrated into the denoiser by simply adjusting the input/output dimensions to accommodate the additional channels. This increase in dimensionality does not overshadow the weights of the original model; instead, only new projection layers are introduced to handle the added modality. To leverage pretrained RGB weights, we can load the pretrained model for the corresponding RGB channels and initialize the remaining weights as zeros. For the conditioning frame, we continue using the first RGB frame and repeat the condition image tensor to prompt the generation of both video and point tracking, as our goal is for the model not to rely on the initial point positions during the inference stage. This strategy is effective because the same reference frame is used for generating the point-tracking data earlier.
@@
In addition to the pixel-wise alignment of video and points during dataset construction, we introduce a cross-attention layer at the input of the latent diffusion model. This layer links the first and second halves of the channel dimensions—corresponding to color information and spatial location in the latent space—to further enhance their alignment. We apply two passes of cross-attention, with each modality alternately serving as the query, while the other provides the key and value in each pass.
@@
Incorporating the aforementioned changes, the training pipeline proceeds as usual: we encode $\mathcal{VP}$ into the latent space $z$, add noise to obtain $z_{t}$ using randomly sampled noise $\epsilon$ and a diffusion step $t$. The denoiser model then predicts the added noise and compares it with the ground-truth noise, where we apply DDIM song2020denoising for diffusion sampling. With this modified pipeline, we fine-tune the diffusion model on PointVid dataset to obtain our video-point generation model. At inference, given an image and a text prompt, the model takes in double-sized random noise and predicts both the RGB video and the corresponding 3D trajectories.
@@ 3.3 Point Regularization
With the model capable of generating pixel-aligned video and points, we leverage this additional modality to enhance RGB video generation. To achieve this, we introduce regularization terms on the point tracking component of the diffusion output, aligning it with the ground truth, which in turn implicitly improves the RGB video component (see Figure 2).
@@
Firstly, we need to recover directly usable point trackings for evaluation in 3D space. Since injecting random noise into the point cloud would likely render it unusable, we employ diffusion sampling to recover a noise-free stage before applying regularization. At each training step, given a latent input $z$ and its corresponding noisy stage $z_{t}$, we apply DDIM denoising to iteratively recover $z_{0}$. Following the approach of yuan2024instructvideo , we adopt gradient checkpointing to freeze gradient flow except during the final step, thereby preventing memory overflow. The denoised latent $z_{0}$ is subsequently decoded into explicit 3D point tracking $\tilde{\mathcal{P}}$, representing the sampled 3D trajectory of the video.
@@
Figure 4: Point Regularization. The reconstructed point cloud in the diffusion output often contains noise and deformations (middle). This issue is mitigated using our point regularization (right). The synthetic point cloud above (e.g., box and shoes falling on the ground) is generated by Kubric  kubric and trained with our pipeline.
@@
With the explicit modeling of trajectories as a 3D representation, we introduce regularization terms to improve the quality of the generated results, focusing on motion fidelity and temporal smoothness. For simplicity, let us assume $\tilde{\mathcal{P}}$ is reshaped to $T\times N\times 3$, and use $\tilde{\mathcal{P}}^{\tau}$ to denote the point cloud of shape $N\times 3$ at time $\tau$.
@@
A common issue that arises during point tracking generation is the presence of excessive noise (see Figure 4 middle). This noise can originate from various sources, such as the variational autoencoder (VAE) used to encode $\mathcal{P}$ into the latent space. High-frequency noise in both spatial and temporal dimensions tends to obscure the main shape and motion, thereby impairing the video model’s ability to track 3D points effectively. To mitigate noise and increase fidelity of reconstruction, we propose the reconstruction loss $\mathcal{L}_{\text{recon}}$ as follows:
$\mathcal{L}_{\text{recon}}=c_{0}\sum_{\tau=0}^{T}\|\tilde{\mathcal{P}}^{\tau}-\mathcal{P}^{\tau}\|+c_{1}\sum_{\tau=1}^{T}\|\tilde{\mathcal{P}}^{\tau}-\tilde{\mathcal{P}}^{\tau-1}\|+c_{2}\sum_{\tau=2}^{T}\|\tilde{\mathcal{P}}^{\tau}-2\tilde{\mathcal{P}}^{\tau-1}+\tilde{\mathcal{P}}^{\tau-2}\|.$ (1)
@@
Here the first term penalizes the difference between generated point cloud and the noise-free ground truth $\mathcal{P}$, ensuring generation fidelity, while the second term penalizes the difference between consecutive time step to encourage temporal smoothness; the third term further enforeces acceleration smoothness, discouraging abrupt changes in velocity. We selected the weights $c_{i}$ so that each term are about the same scale initially.
@@
By combining reconstruction and rigidity loss, we enhance the generation quality of the 3D point cloud, as shown in Figure 4 (right). This improvement in the point space implicitly promotes shape preservation and smooth motion transitions in the RGB video, as our video-point model closely aligns the two modalities. To prevent the geometric regularizations from dominating the semantic aspects, we jointly optimize these regularization terms with the standard diffusion denoiser losses $\mathcal{L}_{\text{diff}}$ song2020denoising during training, resulting in an overall loss of the form (where $\lambda_{[\cdot]}$ are corresponding weights):
$\mathcal{L}=\lambda_{\text{diff}}\mathcal{L}_{\text{diff}}+\lambda_{\text{recon}}\mathcal{L}_{\text{recon}}+\lambda_{\text{rigid}}\mathcal{L}_{\text{rigid}}.$ (3)
@@ 4 Experiments
In this section, we demonstrate the effectiveness of our proposed method through both qualitative and quantitative evaluations. Our fine-tuning approach can be seamlessly integrated into different base video diffusion models. In our experiment, we adopt two baseline image-to-video models with distinct architectures: a UNet-based model, I2VGen-XL zhang2023i2vgen , and a DiT-based model, CogVideoX 1.5 yang2024cogvideox . We refer readers to our supplementary material for more details on the training setup.
Figure 5: Qualitative Comparison. We compare our UNet and DiT models against their respective baselines. The results show that both base models exhibit unrealistic artifacts, such as morphing, while our models ensure smooth transitions in object shape and motion.
@@ 4.1 Qualitative Results
Although our model is designed to generate sequences of 2D images, our point augmentation and regularization improve its ability to capture inherently 3D motions. To evaluate this, we test our models on unseen images and visualize the generated RGB videos. For generation quality, we primarily assess visual plausibility, with a particular focus on shape consistency and preservation. We refer readers to our supplementary document and video for additional results.
Comparison with base models.
We compare our models side by side with their corresponding base models: I2VGen-XL zhang2023i2vgen for the UNet model and CogVideoX 1.5 yang2024cogvideox for the DiT model. Using the same conditioning text and image, we visualize the generated RGB videos in Figure 5. In the videos generated by the base models, noticeable artifacts such as object morphing and unnatural deformations can be observed—for example, the car in (a), the gymnast in (b), and the power saw in (c). In contrast, our models effectively eliminate unnatural shape and appearance distortions, resulting in visually plausible outputs. While our UNet and DiT models differ in architecture and pretrained baselines, we consistently observe a reduction in artifacts compared to their respective base models.
Task-oriented videos.
A prevalent category of real-world videos features humans performing specific tasks (see Figure 1), such as playing with a toy, cooking food, or handling tools. To highlight the challenges of such scenarios, we evaluate three additional video models: SVD (stable-video-diffusion-img2vid-xt) blattmann2023stablevideodiffusionscaling , DynamiCrafter xing2025dynamicrafter , and LTX Video hacohen2024ltx , alongside the two baseline models. We observe that these models often struggle to capture localized details, particularly how human hands interact with objects. This challenge stems from the inherently complex 3D nature of these tasks, which cannot be fully captured using only 2D data. As a result, hands and objects frequently undergo severe distortion or blurring. With our 3D regularization, our model effectively preserves the shapes of both hands and objects, ensuring smooth transitions throughout the video and demonstrating consistent quality improvements over the compared models.
Table 1: Quantitative Evaluation. We evaluate various aspects of our method against other models using the VBench huang2024vbench and VideoPhy bansal2024videophy benchmarks. The evaluated metrics are as follows: (VBench) SC: subject consistency, BC: background consistency, MS: motion smoothness, AQ: aesthetic quality, IQ: imaging quality; (VideoPhy) PC: physical commonsense. Our UNet and DiT models are compared against their respective baselines, with higher scores highlighted in bold, while the highest achiever among all models is underscored.
SVD 0.91807 0.93563 0.97735 0.43276 0.55174 0.34707
LTX Video 0.89001 0.90423 0.99152 0.44670 0.58667 0.33400
I2VGen-XL 0.83704 0.89356 0.96144 0.43140 0.60338 0.32876
Ours (UNet) 0.96003 0.95120 0.98458 0.42246 0.61003 0.37805
CogVideoX 1.5 0.91754 0.92923 0.98798 0.44534 0.60067 0.36549
Ours (DiT) 0.91768 0.92987 0.99201 0.45298 0.60531 0.41384
@@ 4.2 Quantitative Results
To quantitatively evaluate the performance of our model, we adopt a test batch of randomly selected images. The test images are sampled from a single batch of video clips (unseen by our model) from the video dataset chen2024panda , with the initial frame randomly selected within each clip, totaling 372 images. To ensure consistency, we use the same test batch across all quantitative evaluation metrics (including ablation studies in subsection 4.3 and user studies in subsection A.5). In this section, we compare various aspects of video quality between our two models and their respective baseline models, along with two additional models for reference. The results are summarized in Table 1.
First, we assess the general aspects of video generation quality using VBench huang2024vbench , evaluating both temporal quality (subject/background consistency and motion smoothness) and frame-wise quality (aesthetic and imaging quality) of the videos. The results show that our fine-tuned models outperform their baselines in most categories, demonstrating the effectiveness of our approach in improving video generation qualities, particularly in content consistency.
Beyond these general aspects, we are particularly interested in how the injection of 3D awareness enhances alignment with real-world physical principles. It is important to note that determining whether a 2D video is physically accurate is largely subjective, as ground-truth 3D physical states are not available. Therefore, we utilize VideoPhy bansal2024videophy , which employs human annotations to train a discriminator that provides a physical commonsense score. VideoPhy defines physical commonsense from various perspectives, such as solid mechanics, where solid objects are expected to retain their shape and size throughout the video, aligning with our objective. As shown in Table 1, our models achieve a higher score for alignment with physical principles, outperforming their respective baselines and other tested models.
Table 2: Evaluation of Points. We report the MSE of predicted 3D points for the untrained model, the model without regularization, and the model with regularization, respectively.
Item Untrained No Reg With Reg
MSE $\downarrow$ 0.2125 0.0031 0.0020
To further demonstrate the efficacy of our 3D point augmentation and regularization modules, we evaluate the mean squared error (MSE) of the predicted 3D point trajectories during diffusion (using the input point tracking tensor as ground truth), as summarized in Table 2. We observe that, compared to an untrained diffusion model, our network design effectively predicts 3D point trajectories after the augmentation stage and further reduces the error through the regularization stage.
@@ 4.3 Ablation Study
Table 3: Quantitative Ablation Studies. We use VBench huang2024vbench to evaluate the subject and background consistency (SC/BC) of our full model, as well as models trained without point augmentation (PA), without point regularization (PR), cross-attention (CS), and diffusion loss (DL).
Metric Full No PA No PR No CS No DL
SC $\uparrow$ 0.9681 0.9434 0.9544 0.9332 0.7166
BC $\uparrow$ 0.9639 0.9492 0.9522 0.9329 0.9254
We conduct ablation studies on our UNet-based model to evaluate the effectiveness of our training pipeline design. We visualize the inference results in Figure 6 and use VBench huang2024vbench to evaluate the models’ subject and background consistency in Table 3. Our two-stage design of point augmentation and regularization progressively enhances the video model. Compared to training the model with RGB videos only (Figure 6 (i)), our point augmentation alone (Figure 6 (ii)) injects 3D awareness into the model and improves its ability to perceive 3D shapes, as evidenced by more consistent object shapes in the videos. Our point regularization (Figure 6 (iii)) further improves quality by optimizing point generation and implicitly guiding RGB generation towards higher fidelity. Our channel-wise cross-attention mechanism ensures cross-dimensional and cross-modality alignment, without which video quality may degrade (Figure 6 (iv)). Furthermore, in the regularization stage, we adopt a joint training strategy that combines regularization loss with diffusion loss. We observe that dropping diffusion loss (Figure 6 (v)) causes 3D information to dominate semantics, resulting in completely degraded results.
Figure 6: Ablation Studies on Different Components. We compare the results from (iii) our full generation pipeline with (i) training with video only, (ii) training with point augmentation only, (iv) training without channel cross-attention, and (v) training without diffusion loss.
@@ 5 Conclusion and Discussion
To summarize, we propose a generation framework that incorporates 3D-awareness into video diffusion models. By tracking objects in 3D and aligning them in pixel space, we elevate traditional video datasets to a new dimension, revealing out-of-plane information previously inaccessible to video models. Through joint training, the video model learns to perceive 3D shape and motion, acquiring physical commonsense that is inherent in 3D. We then apply regularization to refine the generation process and further enhance the results, eliminating artifacts such as object morphing. Our method can serve as an augmentation for various video models, and we showcase its effectiveness by fine-tuning two widely-used video generation models (I2VGen-XL and CogVideoX 1.5). The results demonstrate a strong capability in reducing non-physical artifacts in video generation, particularly when applied to contact-rich, task-oriented videos.
Motion Magnitude and Smoothness Trade-off.
In the design of our reconstruction loss (Equation 1), we align the generated 3D points with the ground truth positions while enhancing temporal smoothness by minimizing velocity and acceleration. While we acknowledge that this loss introduces a trade-off between motion magnitude and smoothness, the global dynamics are primarily governed by the position term (to which we assign larger weights). As a result, we observe faithful reconstruction of ground-truth dynamics during training (see Table 2). Moreover, discouraging abrupt changes in velocity is often beneficial, as they typically induce morphing artifacts in generated video (see Figure 5). We refer the reader to subsection A.2 for more details on the training setup.
@@
We acknowledge a few limitations of our framework. The degree of 3D-awareness, rooted in our video-point joint training, is constrained by the resolution of our 3D points. While our method does not require highly accurate point tracking for 3D prior, objects that are not sufficiently represented may lead to suboptimal 3D perception by the video model. Recent advancements in 3D point tracking ngo2024delta ; xiao2025spatialtrackerv2 could be deployed to enhance the 3D perception in our model. Additionally, the generation quality of our model is limited by the base diffusion backbone and could potentially be improved by employing larger-scale models. We leave these aspects as directions for future exploration.
Concurrent Works.
Concurrently, several studies have emerged to enhance shape and motion consistency in video generation. chefer2025videojam uses optical flow as a representation of object motion and does not explicitly model object geometry. jeong2025track4gen proposes using point tracking to improve coherence; however, it only utilizes 2D pixel correspondence without 3D spatial information. On the contrary, our method explicitly models the 3-dimensional geometry and motion to provide spatial guidance to the diffusion model.
Appendix A Technical Appendices and Supplementary Material
Supplementary to the main paper, we provide more details on dataset creation in subsection A.1, more details on implementation in subsection A.2, additional qualitative results in subsection A.3, and a human preference evaluation in subsection A.5.
A.1 PointVid Dataset Generation
In this section, we present additional details on generating our 3D-aware video dataset, PointVid. A visualization of the generation pipeline is provided in the main paper. Our pipeline requires no additional annotations beyond the video-text pairs in common video datasets.